# A Survey of Deep Learning for Complex Speech Spectrograms

### v1

@@ 5 Training Strategies and Loss Functions
When DNNs are used to solve problems related to complex spectrograms, the main goal is to estimate the target complex spectrograms from the input. The estimation approaches vary, and a range of different training strategies have been employed. These strategies can generally be categorized into direct estimation (mapping) and indirect estimation (masking). Furthermore, loss functions are used to measure how closely the algorithm’s output matches the target. Given that complex spectrograms can be represented in various coordinate systems, multiple loss functions have been explored. The first subsection below will provide an overview of the training strategies employed in complex spectrogram-related works. The second subsection will summarize the loss functions utilized for learning complex spectrogram-related targets.
5.1 Training Strategies
@@
Moreover, recent research has witnessed the emergence of integrated approaches combining masking and mapping techniques, particularly in the context of magnitude and phase spectrogram processing (Yin et al. (2020); Lu et al. (2023); Kühne et al. (2025)). Building on the success in magnitude spectrogram processing, masking remains central in these frameworks — responsible for magnitude estimation, while a parallel branch handles phase mapping.
@@ 5.2 Loss Functions
An well-designed loss function can substantially improve model performance without increasing computational overhead, thereby serving as a critical component in the training process. Given that complex spectrograms can be represented in multiple coordinate systems and subsequently transformed back to the time domain, a variety of loss functions have been proposed to align with these distinct representations. In this survey, we systematically present the loss functions from four perspectives: phase-related loss functions, complex-domain loss functions, time-domain loss functions, and multi-domain loss functions. The discussion of magnitude-domain loss functions is omitted, as they have been extensively studied and this survey focuses on complex spectrogram-realted works.
@@ 5.2.1 Phase-Related Loss Functions
Estimating phase spectrograms is a challenging task due to their intrinsic properties, such as the lack of distinct structure and the complexities introduced by phase wrapping. Research related to phase has been explored in various domains, including phase retrieval (Takamichi et al. (2018)), speech enhancement (Yin et al. (2020)), and speech separation (Wang et al. (2019)). Details about the natural properties of phase and recent advances in phase retrieval will be discussed in Subsection 6.1. In the following, we outline the phase-related loss functions that are frequently employed.
@@
Here, $\bm{\theta}$ represents a periodic variant, which may refer not only to the phase itself but also to its derivatives. The indices $k$ and $l$ correspond to the frequency bin and time frame, respectively. Furthermore, Thieling et al. (2021) proposed a regularized cosine loss, as:
$\mathcal{L}_{\text{reg}}(\bm{\theta},\hat{\bm{\theta}})=\sum_{l,k}\{-\cos(\theta_{l,k}-\hat{\theta}_{l,k})+\frac{(\theta_{l,k}-\hat{\theta}_{l,k})^{4}}{4000}\}.$ (39)
Thieling et al. (2021) demonstrated that regularized cosine loss enhances training stability and provides a slight improvement in group delay estimation.
@@ 5.2.4 Multi-Domain Loss Functions
Despite the intuitive appeal of loss functions in the complex or time domains, experimental results have shown only limited performance improvements, particularly in terms of magnitude-based metrics (Pandey and Wang (2019); Wang et al. (2021b)). Consequently, several studies have explored the design of loss functions with the aim of improving the model performance on evaluation metrics. Speech signal evaluation typically takes multiple factors into account. For instance, SSNR and SI-SDR measures are computed in the time domain, whereas perceptual evaluation of speech quality (PESQ), short-time objective intelligibility (STOI), extended STOI (ESTOI), and log-spectral distortion (LSD) primarily focus on the magnitude spectrograms. Since evaluation metrics account for multiple domains, a feasible approach to loss function design is to incorporate losses from multiple domains, ensuring a more comprehensive optimization strategy.
@@
Wang and Wang (2020) further advanced the field by proposing a loss function that applies the L1 loss to the real, imaginary and magnitude spectrograms. Its effectiveness has been validated in multiple domains, including speech resynthesis (Xie et al. (2024)), speech enhancement (Wang et al. (2022)), and speech separation (Wang et al. (2021b)). The promising performance improvements achieved by this multi-domain loss functions have spurred further research into loss function design. Several subsequent studies have explored this topic, including theoretical analyses (Wang et al. (2021b)) and experimental comparisons (Braun and Tashev (2021)) to gain deeper insight into effectiveness. For instance, Wang et al. (2021b) provided a theoretically analysis of the performance improvement, framing it as a compensation problem. Let $S(l,k)$ and $\hat{S}(l,k)$ denote the ground truth and the estimated spectrogram bin, respectively. In general, phase estimation is challenging, meaning that $\angle\hat{S}(l,k)$ can deviate significantly from $\angle{S}(l,k)$, especially when the input signal has a low SNR. If the phase error $\angle S(l,k)-\angle\hat{S}(l,k)$ exceeds $\pi/2$, loss functions that consider only the direct distance $S(l,k)-\hat{S}(l,k)$ may drive the network to predict near-zero magnitudes, i.e. $|\hat{S}(l,k)|$ approaching zero, ultimately leading to ineffective predictions. This explains why training solely with phase-aware losses alone does not yield strong performance on magnitude-sensitive metrics such as PESQ and ESTOI. Incorporating a magnitude loss into the total loss function helps mitigate this issue by encouraging the DNN to achieve a balanced optimization in both complex and magnitude domains. Moreover, Braun and Tashev (2021) investigated the impact of various loss functions on evaluation performance in a lightweight RNN architecture for online speech enhancement. Their study demonstrated that combining magnitude-only and phase-aware losses consistently led to performance improvements. Additionally, they found that using compressed spectrals as features can further enhance performance, which is consistent with the results from Li et al. (2021).
@@ 6 Applications
This chapter explores three major applications in which DNN-based approaches have shown substantial impact on tasks involving complex spectrograms: (1) phase retrieval (also known as phase reconstruction), which aims to estimate the phase spectrogram from a given magnitude spectrogram; (2) speech enhancement, focused on improving the clarity, intelligibility, and overall quality of speech signals in the presence of noise, reverberation, or other distortions; and (3) speech separation, which seeks to isolate individual speech signals from mixtures containing overlapping voices or background noise.
@@ 6.1 Phase Retrieval
Nonetheless, direct phase retrieval using DNNs remains a challenging task. One major obstacle is the phase wrapping issue. As phase is naturally periodic, its value is only available between $[-\pi,\pi)$. This nature not only causes discontinuity in training but also renders typical regression loss like MSE ineffecient. Another significant challenge lies in phase sensitivity to waveform shifts. Unlike the magnitude spectrogram, phase is highly sensitive to time shifts, with even minor shifts causing substantial variations. Moreover, additional complexities such as sign indertermination (Sturmel et al. (2011)) further complicate the phase retrieval process, making it a nontrival problem for deep learning-based approaches.
@@
One approach is to directly estimate the complex spectrogram, thus bypassing the need for direct phase estimation. For instance, Oyamada et al. (2018) proposed using GAN to reconstruct the complex spectrogram from the known magnitude. During training, the generator reconstructs the complex spectrogram, while the discriminator evaluates it against the ground truth. Ai and Ling (2023) proposed a discriminative framework for phase retrieval via parallel spectrogram estimation. Their architecture employs a dual-path mechanism that simultaneously estimates pseudo-real and pseudo-imaginary components to facilitate phase reconstruction.
@@ 6.2 Speech Enhancement
In this subsection, we present a review of recent advancements in processing complex spectrograms for speech enhancement. While previous studies have categorized speech enhancement methodologies into mapping and masking approaches based on their training strategies (Wang and Chen (2018); Michelsanti et al. (2021)), we adopt a novel classification framework that categorizes these works into two paradigms: discriminative models and generative models. This shift is motivated by the growing attention that generative models have recently received in the field of speech enhancement.
6.2.1 Discriminative Methods
@@
With the growing interest in complex-valued neural networks (Trabelsi et al. (2018)) and the fact that spectrograms are naturally complex-valued, several studies have extended DNNs from the real-valued domain to the complex domain. For instance, Choi et al. (2019) proposed deep complex U-Net (DCUNET). Its components, including complex-valued convolutional layers, complex batch normalization and leaky $\mathbb{C}$ReLU, have been discussed in Sections 3 and 4. With incorporating a time-domain loss function during training, DCUNET was subsequently adopted for speech enhancement through complex mask estimation. It is worth noting that experimental results from Choi et al. (2019) reveal that, consistent with observations in real-valued neural networks, a deeper complex-valued architecture exhibits superior enhancement performance.
@@
While aforementioned CNN-based models have shown promising results, their limited receptive fields restrict them to local feature extraction. However, speech signals require modeling long-range dependencies across time and frequency. To overcome this, recent work has explicitly incorporated joint time-frequency context modeling. Tan and Wang (2019a, b) introduced a real-valued convolutional recurrent network (CRN) specifically designed for complex spectrogram mapping. Experimental results demonstrate the effectiveness of combining the strengths of convolutional operations for local feature extraction and recurrent structures for temporal modeling. Furthermore, Hu et al. (2020) proposed a complex-valued architecture known as the Deep Complex Convolution Recurrent Network (DCCRN). By estimating a complex-valued mask and optimizing with an SI-SDR loss, DCCRN achieved superior performance over existing baselines while maintaining significantly lower computational complexity. Recent advancements in neural networks have spurred a wave of relevant research efforts. Modern techniques such as densely connected networks (Huang et al. (2017)), dual-path RNNs (Luo et al. (2020)), attention mechanisms (Vaswani et al. (2017)) have been employed to improve the enhanced performance (Tan et al. (2021); Le et al. (2021); Zhao et al. (2021)).
While phase spectrogram modeling remains challenging, several studies have pursued speech enhancement via joint magnitude-phase processing (Yin et al. (2020); Yu et al. (2022); Fu et al. (2022); Lu et al. (2023); Kühne et al. (2025)). These approaches often adopt strategies for handling the phase component that are either inspired by or closely related to phase retrieval techniques. The architectural design of such models typically employs a dual-path  [^1] Two distinct ”dual-path” mechanisms have been proposed in literatures. The first is exemplified by the dual-path RNN, which incorporates intra- and inter-chunk processing to capture both local and global information. Representive works include dual-path RNN (Luo et al. (2020)), TF-GridNet (Wang et al. (2023)) and so on. The second dual-path mechanism, discussed here, employs a parallel architecture, where magnitude and phase information are processed in different branches. Thus, this is also called as dual-branch in some works. mechanism, where one branch is dedicated to processing the magnitude spectrogram, while the other focuses on either the phase or the complex spectrogram. Crucially, these branches are interconnected through an information exchange mechanism, which serves two primary purposes: facilitating more effective phase modeling and boosting the overall performance of the system. This cross-branch interaction has proven to be essential in achieving better speech enhancement results in the dual-path mechanism (Yin et al. (2020)).
6.2.2 Generative Model-based Approaches
Unlike discriminative methods that directly learn a mapping from noisy to clean speech, generative models can learn to generate realistic and high-quality speech by capturing the underlying structure of clean speech data. This allows them to effectively suppress noise while preserving speech naturalness and intelligibility.
Although most early studies on GAN-based speech enhancement focused primarily on the magnitude domain (Michelsanti and Tan (2017)) or the time domain (Pascual et al. (2017)), recent research has extended the application of GANs to the complex spectrogram domain (Kothapally and Hansen (2022); Cao et al. (2022); Zhang et al. (2024); Lu et al. (2023); Kühne et al. (2025)). In these studies, the generator networks — often inspired by advancements in discriminative modeling — are trained to estimate either masks or direct mappings for complex spectrograms (Kothapally and Hansen (2022); Cao et al. (2022)). To further enhance model performance, the discriminators are designed to differentiate between enhanced and clean speech, employing various optimization strategies during training (Kothapally and Hansen (2022); Cao et al. (2022); Zhang et al. (2024)). Experimental results demonstrate that adversarial training contributes to improved enhancement performance (Cao et al. (2022)), further validating the effectiveness of applying GANs to speech enhancement tasks. Moreover, to address the challenge of phase enhancement, some studies have incorporated loss functions inspired by phase retrieval techniques, leading to additional performance gains (Zhang et al. (2024)).
In speech enhancement on complex spectrograms, several diffusion model-based approaches have been explored. Richter et al. (2023) proposed one approach which leverages continuous SDEs to model a progressive transformation of the mean from clean speech to noisy speech, enabling high-quality enhancement by reversing this diffusion process. Besides, some works combine diffusion models with discriminative models, where the discriminative model provides an initial estimate of clean speech, and the diffusion model refines it by further reducing distortions through techniques like denoising diffusion restoration and stochastic regeneration (Wang and Wang (2023); Sawata et al. (2023); Shi et al. (2024); Lemercier et al. (2023a)). This hybrid method avoids generating artifacts such as vocalizing and breathing effects while significantly reducing computational costs (Lemercier et al. (2023a)). Moreover, advanced unconditional diffusion models have been applied to address the inverse problem and have further been utilized for speech dereverberation in the work from Lemercier et al. (2023b). For an in-depth discussion of diffusion models and complex spectrogram processing, please refer to Section 7 and recent work from Lemercier et al. (2024).
6.3 Speech Separation

### v2

@@ 6 Training Strategies and Loss Functions
For real-time speech applications such as hearing aids and telecommunication systems, low latency is essential. Conventional spectrogram-based methods often require long window length in STFT calculation for better performance, but at the cost of additional algorithm latency. To address this, approaches such as asymmetric windowing and future-frame prediction have been explored. In addition, when DNNs are used to solve problems related to complex spectrograms, the main goal is to estimate the target complex spectrograms from the input. The estimation approaches vary, and a range of different training strategies have been employed. These strategies can generally be categorized into direct estimation (mapping) and indirect estimation (masking). Finally, loss functions are employed to measure how closely the algorithm’s output matches the target. Since complex spectrograms can be represented in various coordinate systems, multiple loss functions have been proposed and studied. The first subsection will summarize the low-latency techniques that enable practical deployment. The second subsection below will provide an overview of the training strategies employed in complex spectrogram-related works. The third subsection will discuss the loss functions utilized for learning complex spectrogram-related targets.
6.1 Low-latency
Low-latency is a critical concern, especially in real-time speech processing systems where preserving perceptual quality is essential. Here latency refers to the time delay between the input of a speech signal and the production of the processed output. Latency generally consists of two components: algorithmic latency and hardware-induced computational latency (Dementyev et al. (2025); Wang et al. (2023b); Zheng et al. (2022)).
The STFT window length largely determines the algorithmic latency in complex spectrogram–based speech processing. Owing to the inherent trade-off between time and frequency resolution, the choice of window length directly affects model performance. Shorter windows reduce latency but yield poor frequency resolution, thereby limiting the capability of DNN models. Conversely, longer windows provide higher frequency resolution and richer spectral details, which further enhance the modeling capacity of DNN frameworks, but often fail to satisfy the strict latency requirements of real-time applications (Wu and Braun (2025); Wang et al. (2023b)).
To enable real-time processing with high speech quality, several strategies have been developed for handling complex spectrograms (Wu and Braun (2025)). One approach is to use asymmetric STFT windowing (Mauler and Martin (2007); Wang et al. (2021a, 2023b)). Here, a larger analysis window is used to compute the input STFT features, while a shorter synthesis window is employed for the ISTFT. This approach has been shown to reduce latency in complex spectrogram-based models to as low as 4 ms (Wang et al. (2023b)).
Latency can also be reduced through future-frame prediction (Iotov et al. (2022); Wang et al. (2023b)). This technique predicts the features of the next frame directly from past and current observations, eliminating the need to wait for the actual future frame. With a 50% overlap in STFT calculation, the total latency can be effectively reduced to the hop size. Moreover, for a given latency constraint, future-frame prediction allows the use of a longer analysis window, thereby improving the frequency resolution of the input spectrograms.
6.2 Mapping and Masking
@@
Moreover, recent research has witnessed the emergence of integrated approaches combining masking and mapping techniques, particularly in the context of magnitude and phase spectrogram processing (Yin et al. (2020); Lu et al. (2023); Kühne et al. (2025b)). Building on the success in magnitude spectrogram processing, masking remains central in these frameworks — responsible for magnitude estimation, while a parallel branch handles phase mapping.
@@ 6.3 Loss Functions
A well-designed loss function can substantially improve model performance without increasing computational overhead, thereby serving as a critical component in the training process. Given that complex spectrograms can be represented in multiple coordinate systems and subsequently transformed back to the time domain, a variety of loss functions have been proposed to align with these distinct representations. In this survey, we systematically present the loss functions from four perspectives: phase-related loss functions, complex-domain loss functions, time-domain loss functions, and multi-domain loss functions. The discussion of magnitude-domain loss functions is omitted, as they have been extensively studied and this survey focuses on complex spectrogram-realted works.
@@ 6.3.1 Phase-Related Loss Functions
Estimating phase spectrograms is a challenging task due to their intrinsic properties, such as the lack of distinct structure and the complexities introduced by phase wrapping. Research related to phase has been explored in various domains, including phase retrieval (Takamichi et al. (2018)), speech enhancement (Yin et al. (2020)), and speaker separation (Wang et al. (2019)). Details about the natural properties of phase and recent advances in phase retrieval will be discussed in Subsection 7.1. In the following, we outline the phase-related loss functions that are frequently employed.
@@
Here, $\bm{\theta}$ represents a periodic variant, which may refer not only to the phase itself but also to its derivatives. The indices $k$ and $l$ correspond to the frequency bin and time frame, respectively. Furthermore, Thieling et al. (2021) proposed a regularized cosine loss which yields more stable training.
@@ 6.3.4 Multi-Domain Loss Functions
Despite the intuitive appeal of loss functions in the complex or time domains, experimental results have shown only limited performance improvements, particularly in terms of magnitude-based metrics (Pandey and Wang (2019); Wang et al. (2021c)). Consequently, several studies have explored the design of loss functions with the aim of improving the model performance on evaluation metrics. Speech signal evaluation typically takes multiple factors into account. For instance, SSNR and SI-SDR measures are computed in the time domain, whereas perceptual evaluation of speech quality (PESQ), short-time objective intelligibility (STOI), extended STOI (ESTOI), and log-spectral distortion (LSD) primarily focus on the magnitude spectrograms. Since evaluation metrics account for multiple domains, a feasible approach to loss function design is to incorporate losses from multiple domains, ensuring a more comprehensive optimization strategy.
@@
Wang and Wang (2020) further advanced the field by proposing a loss function that applies the L1 loss to the real, imaginary and magnitude spectrograms. Its effectiveness has been validated in multiple domains, including speech resynthesis (Xie et al. (2024)), speech enhancement (Wang et al. (2022)), and speaker separation (Wang et al. (2021c)). The promising performance improvements achieved by this multi-domain loss functions have spurred further research into loss function design. Several subsequent studies have explored this topic, including theoretical analyses (Wang et al. (2021c)) and experimental comparisons (Braun and Tashev (2021)) to gain deeper insight into effectiveness. For instance, Wang et al. (2021c) provided a theoretical analysis of the performance improvement, framing it as a compensation problem. Let $S(l,k)$ and $\hat{S}(l,k)$ denote the ground truth and the estimated spectrogram bin, respectively. In general, phase estimation is challenging, meaning that $\angle\hat{S}(l,k)$ can deviate significantly from $\angle{S}(l,k)$, especially when the input signal has a low SNR. If the phase error $\angle S(l,k)-\angle\hat{S}(l,k)$ exceeds $\pi/2$, loss functions that consider only the direct distance $S(l,k)-\hat{S}(l,k)$ may drive the network to predict near-zero magnitudes, i.e. $|\hat{S}(l,k)|$ approaching zero, ultimately leading to ineffective predictions. This explains why training solely with phase-aware losses alone does not yield strong performance on magnitude-sensitive metrics such as PESQ and ESTOI. Incorporating a magnitude loss into the total loss function helps mitigate this issue by encouraging the DNN to achieve a balanced optimization in both complex and magnitude domains. Moreover, Braun and Tashev (2021) investigated the impact of various loss functions on evaluation performance in a lightweight RNN architecture for online speech enhancement. Their study demonstrated that combining magnitude-only and phase-aware losses consistently led to performance improvements. Additionally, they found that using compressed spectrals as features can further enhance performance, which is consistent with the results from Li et al. (2021). Subsequently, the design of loss functions has been extended to incorporate penalties in multiple domains. For example, Wang and Wang (2021) proposed an objective function under a cascade architecture, which combines mask-based, complex-domain, and time-domain components.
Futhermore, most of the aforementioned methods compute the loss function using a single set of STFT parameters, thereby limiting the optimization of complex spectrogram to a single resolution. To overcome this limitation, Yu et al. (2022b) recently proposed a multi-resolution loss function that combines losses computed across multiple STFT window lengths into a unified objective. Experimental results have shown that incorporating penalties computed over multiple window lengths can further improve model performance.
@@ 7 Applications
Complex spectrogram processing has been applied to a variety of speech tasks, including but not limited to phase retrieval, speech enhancement, speaker separation, emotion recognition (Guo et al. (2025)) and sound source localization (Kim et al. (2024)). However, it should be noted that not all speech-related tasks require explicit phase modeling, since many of them can achieve satisfactory performance using magnitude features alone. In speech signal processing, the primary benefits of incorporating phase information are often associated with improving perceptual quality for human listeners.
In this chapter, we focus on three major applications where DNN-based approaches have demonstrated substantial advancements in leveraging complex spectrograms: (1) phase retrieval (also known as phase reconstruction), which aims to estimate the phase spectrogram from a given magnitude spectrogram; (2) speech enhancement, focused on improving the clarity, intelligibility, and overall quality of speech signals in the presence of noise, reverberation, or other distortions; and (3) speaker separation, which seeks to isolate individual speech signals from mixtures containing overlapping voices.
@@ 7.1 Phase Retrieval
Nonetheless, direct phase retrieval using DNNs remains a challenging task. One major obstacle is the phase wrapping issue. As phase is naturally periodic, its value is only available between $\pi,\pi)$. This nature not only causes discontinuity in training but also renders typical regression loss like MSE ineffecient. Another significant challenge lies in phase sensitivity to waveform shifts. Unlike the magnitude spectrogram, phase is highly sensitive to time shifts, with even minor shifts causing substantial variations. Moreover, additional complexities such as sign indertermination (Sturmel et al. (2011)) further complicate the phase retrieval process, making it a nontrival problem for deep learning-based approaches.
@@
One approach is to directly estimate the complex spectrogram, thus bypassing the need for direct phase estimation. For instance, Oyamada et al. (2018) proposed using GAN to reconstruct the complex spectrogram from the known magnitude. During training, the generator reconstructs the complex spectrogram, while the discriminator evaluates it against the ground truth. Ai and Ling (2023) proposed a discriminative framework for phase retrieval via parallel spectrogram estimation. Their architecture employs a dual-branch mechanism that simultaneously estimates pseudo-real and pseudo-imaginary components to facilitate phase reconstruction.
@@ 7.2 Speech Enhancement
In this subsection, we review recent advancements in processing complex spectrograms for speech enhancement, especially focusing on discriminative models. For generative model-based speech enhancement methods, please referred to Section 8.
@@
With the growing interest in CVNNs (Trabelsi et al. (2018)) and the fact that spectrograms are naturally complex-valued, several studies have extended DNNs from the real-valued domain to the complex domain. For instance, Choi et al. (2019) proposed deep complex U-Net (DCUNET). Its components, including complex-valued convolutional layers, complex batch normalization and leaky $\mathbb{C}$ReLU, have been discussed in Sections 3 and 4. With incorporating a time-domain loss function during training, DCUNET was subsequently adopted for speech enhancement through complex mask estimation. It is worth noting that experimental results from Choi et al. (2019) reveal that, consistent with observations in RVNNs, a deeper complex-valued architecture exhibits superior enhancement performance.
@@
While aforementioned CNN-based models have shown promising results, their limited receptive fields restrict them to local feature extraction. However, speech signals require modeling long-range dependencies across time and frequency. To overcome this, recent work has explicitly incorporated joint time-frequency context modeling. Tan and Wang (2019a, b) introduced a real-valued convolutional recurrent network (CRN) specifically designed for complex spectrogram mapping. Experimental results demonstrate the effectiveness of combining the strengths of convolutional operations for local feature extraction and recurrent structures for temporal modeling. Furthermore, Hu et al. (2020) proposed a complex-valued architecture known as the Deep Complex Convolution Recurrent Network (DCCRN). By estimating a complex-valued mask and optimizing with an SI-SDR loss, DCCRN achieved superior performance over existing baselines while maintaining significantly lower computational complexity.
Subsequent research has focused on leveraging both cross-band and sub-band information to more effectively capture time–frequency dependencies. This joint modeling strategy allows the network to exploit fine-grained spectral details within sub-bands while also accounting for global interactions across bands, which has been shown to enhance both objective metrics and perceptual quality in enhancement. Representative real-valued models in this direction include FullSubNet (Hao et al. (2021)), DPT-FSNet (Dang et al. (2022)), CMGAN (Cao et al. (2022)), and BSRNN (Yu and Luo (2023)). Similar ideas have also been explored in complex-valued models, such as DCCRN+ (Lv et al. (2021)).
Although modeling the phase spectrogram remains a significant challenge, recent studies have investigated speech enhancement through joint magnitude–phase processing (Yin et al. (2020); Yu et al. (2022a); Fu et al. (2022); Lu et al. (2023); Kühne et al. (2025b)). As discussed in Section 5.2.2, their architectures typically adopt either a dual-branch mechanism—where one branch estimates the magnitude spectrogram and the other focuses on phase or complex representations with cross-branch information exchange—or an encoder–decoder structure, and are therefore not elaborated here. Building on the success of magnitude spectrogram processing, masking remains a central component for magnitude estimation in these works, while a parallel branch is often employed for phase mapping.
7.3 Speaker Separation